# Electronic modes induced by spin and charge perturbations in Mott and Kondo insulators

### v1

@@ VI.3 Spin-charge separation of interacting metals on a chain
Spin-charge separation means that the lowest spin-excitation energy differs from the lowest charge-excitation energy. In interacting metals on a chain, it is known that the velocity of spin excitation generally differs from that of charge excitation; spin-charge separation occurs in the energy scale of $\mathcal{O}(\frac{1}{L})$, where $L$ denotes the number of sites on a chain Essler ; HaldaneTLL ; TomonagaTLL ; LuttingerTLL ; MattisLiebTLL ; Giamarchi . This characteristic is contrasted with that of a Fermi liquid LandauFL , where spin and charge excitations are obtained as particle-hole excitations of fermionic quasiparticles. In a Fermi liquid, the lowest spin-excitation energy is the same as the lowest charge-excitation energy, and spin-charge separation does not occur.
@@ VI.4 Spin-charge separation of strongly correlated insulators
VI.5 Relation between doping-induced ingap electronic modes and spin-charge separation of strongly correlated insulators
According to the quantum-number analysis (selection rules) for doping-induced modes in Sec. IV.2, the dispersion relations of the doping-induced ingap modes are expressed [Eqs. (43) and (48)] as
@@
which implies that the emergent ingap electronic modes exhibit the spin-excitation dispersion relation ($\pm e^{\rm spin}_{\pm{\bm{k}}}$) shifted by the Fermi momentum ($\mp{\bm{k}}_{\rm F}^{\mp}$) from the band edges ($\mu_{\mp}$).
@@
Thus, the doping-induced ingap modes reflect spin-charge separation of strongly correlated insulators KohnoMottT ; KohnoTinduced ; KohnoRPP ; Kohno1DHub ; Kohno2DHub ; Kohno1DtJ ; Kohno2DtJ ; KohnoDIS ; KohnoHubLadder ; KohnoSpin ; KohnoAF ; KohnoGW ; KohnoKLM ; KohnottpHub ; KohnottpJ ; these do not appear in conventional band insulators without spin-charge separation.
The emergence of ingap modes in the electronic excitation from the lowest-energy state of a nonzero value of $\eta$ at half filling ($\eta^{z}=0$) [Sec. IV.6; Figs. 6(c), 6(f), 7(c), 8(c), 8(f), 8(i), 8(l), 8(o), 8(r), 8(u), 8(x), 9(c), 9(f), 9(i), and 9(l)] also reflects spin-charge separation.
VI.6 Relation between magnetization-induced ingap electronic modes and spin-charge separation of strongly correlated insulators
@@ VI.6.1 Emergence of magnetization-induced modes within the band gap in spin-gapped Mott and Kondo insulators
Similarly, according to the quantum-number analysis (selection rules) for magnetization-induced modes in Sec. V.2, the dispersion relations of the magnetization-induced ingap modes are expressed [Eqs. (78) and (79)] as
@@
The spin gap at $H=0$ is denoted as $\Delta_{\rm spin}$.
@@
From the above consideration, the formation of the spin mode (triplon mode) separated from the particle-hole continuum can be considered the key characteristic for the emergence of electronic modes with magnetization. As shown in Figs. 4 and 5, the spin mode (triplon mode) appears below the particle-hole continuum, and charge mode appears above the particle-hole continuum in spin-gapped Mott and Kondo insulators. Because the lowest excitation energy of the spin mode is lower than that of the particle-hole continuum, which is equal to the lowest charge-excitation energy, the formation of the spin mode below the particle-hole continuum is a manifestation of spin-charge separation in spin-gapped Mott and Kondo insulators.
Thus, the emergence of magnetization-induced modes within the band bap reflects spin-charge separation, which is realized as the formation of the spin mode below the particle-hole continuum, in spin-gapped Mott and Kondo insulators.
The emergence of ingap modes in the electronic excitation from the lowest-energy state of a nonzero value of $S$ without magnetization at half filling [Sec. V.6; Figs. 11(c), 11(f), 11(i), 11(l), 11(o), 11(r), 11(u), 11(x), 12(c), 12(f), 12(i), and 12(l)] also reflects spin-charge separation.
@@ VI.6.4 Strongly correlated insulators of attractively interacting electrons
In the insulators of attractively interacting electrons ($U<0$) which are obtained using the Shiba transformation (Sec. V.5) from the repulsive case ($U>0$), because the roles of spin and charge are interchanged by the Shiba transformation [Eqs. (102) and (103)], the charge mode appears below the particle-hole continuum. In this case, the doping-induced modes appear within the band gap, reflecting spin-charge separation (the lowest charge-excitation energy is lower than the lowest spin-excitation energy, which is opposite to the repulsive case).
@@ VI.7 Implications
Although various interpretations of electronic states in strongly correlated systems have been considered KohnoMottT ; KohnoTinduced ; KohnoRPP ; Kohno1DHub ; Kohno2DHub ; Kohno1DtJ ; Kohno2DtJ ; KohnoDIS ; KohnoHubLadder ; KohnoSpin ; KohnoAF ; KohnoGW ; KohnoKLM ; KohnottpHub ; KohnottpJ ; Eskes ; DagottoDOS ; CPTPRL ; CPTPRB ; CPTHanke ; TsunetsuguRMP ; localRungHubLadder ; lightPump1dHub ; pumpProbe1dEHM ; photoSubbands ; FKFloquetDMFT ; noneqDMFTRMP ; DCAkw1dHub ; entropyCoolingDMFT ; etaPairingSuper ; nonthermalSuper ; KanekoEtaPairing ; etaPairingHubLad ; Essler ; HaldaneTLL ; TomonagaTLL ; LuttingerTLL ; MattisLiebTLL ; Giamarchi ; DDMRGAkw ; HubbardI ; LandauFL ; BrinkmanRice ; SlaterAF ; PennAF ; ImadaRMP ; SakaiImadaPRL ; SakaiImadaPRB ; SakaiJPSJ ; ImadaCofermionPRL ; ImadaCofermionPRB ; PhillipsRMP ; PhillipsRPP ; EderOhta2DHub ; EderOhtaIPES ; PreussQP ; GroberQMC ; NoceraDMRT_T ; Akw2dtJ_T , the view presented in this paper along with Refs. [KohnoTinduced, ; KohnoMottT, ; KohnoRPP, ; Kohno1DHub, ; Kohno2DHub, ; Kohno1DtJ, ; Kohno2DtJ, ; KohnoDIS, ; KohnoHubLadder, ; KohnoSpin, ; KohnoAF, ; KohnoGW, ; KohnoKLM, ; KohnottpHub, ; KohnottpJ, ] can consistently explain unconventional features of the electronic states induced not only by temperature increase, doping, and magnetization but also by spin and charge perturbations that cause nonequilibrium states around Mott and Kondo insulators (Secs. IV and V) and can clarify the differences from the conventional features of the temperature-, doping-, and magnetization-independent band structure of noninteracting band insulators (Sec. VI).
According to this view, electronic modes, whose dispersion relations can be expressed as combinations of the electronic dispersion relation and spin or charge dispersion relation of unperturbed systems, can be induced by various spin and charge perturbations as well as by temperature increase in strongly correlated insulators KohnoMottT ; KohnoTinduced ; KohnoRPP ; Kohno1DHub ; Kohno2DHub ; KohnoDIS ; Kohno1DtJ ; Kohno2DtJ ; KohnoSpin ; KohnoAF ; KohnoGW ; KohnottpHub ; KohnottpJ ; KohnoHubLadder ; KohnoKLM . This characteristic can be applied to band-structure engineering, which makes it possible to increase the number of electronic bands, control the band gap, and even change the insulating band structure to metallic at half filling by appropriately tuning and selecting the perturbations and strongly correlated materials.
@@ VII Summary
This study demonstrated that various spin and charge perturbations can induce electronic modes and change the band structure in strongly correlated insulators, provided that a macroscopic number $\mathcal{O}(N_{\rm u})$ of spins or charges are excited. The origins and dispersion relations of the emergent electronic modes were elucidated by investigating the selection rules and using the Bethe ansatz and effective theory for weak inter-unit-cell hopping. The theoretical results were validated by numerical calculations for the 1D and 2D HMs, PAMs, KLMs, and ladder and bilayer HMs.
@@
Regarding charge perturbations, not only the chemical-potential shift that causes hole or electron doping but also the perturbation that generates the lowest-energy state of a nonzero value ($\bar{\eta}$}) of $\eta$ at half filling ($N_{\rm h}=0$) can induce electronic modes. The induced electronic modes at half filling exhibit the same dispersion relations as those of the doping-induced modes at $|N_{\rm h}|=2{\bar{\eta}}$ for $\mu=0$, reflecting $\eta$-SU(2) symmetry. In addition, the charge fluctuation that generates effectively a single dominant state at a momentum in the charge excitation can induce electronic modes. The charge perturbations at half filling can be realized by photon (light) radiation.
@@
Regarding spin perturbations, the magnetic field that magnetizes spin-gapped Mott and Kondo insulators can induce electronic modes within the band gap. The spin perturbation that generates the lowest-energy state of a nonzero value ($\bar{S}$}) of $S$ without magnetization ($S^{z}=0$) can also induce electronic modes, which exhibit the same dispersion relations as those of the magnetization-induced modes at $|S^{z}|={\bar{S}}$ for $H=0$, reflecting spin-SU(2) symmetry. In addition, the spin fluctuation that generates effectively a single dominant state at a momentum in the spin excitation can induce electronic modes. The spin perturbations without a magnetic field can be realized through neutron or photon radiation.
@@
The dispersion relations of the emergent electronic modes can be simply expressed as combinations of the electronic dispersion relation and spin or charge dispersion relation of unperturbed systems. The results can be used to interpret changes in the band structure caused by perturbations, including the emergence of electronic states in nonequilibrium systems. Furthermore, the energy and momentum regimes of the emergent electronic states can be controlled by tuning and selecting the perturbations and strongly correlated materials based on the present results. This approach can be applied to the basic principle of band-structure engineering for strong-correlation electronics, utilizing the unconventional characteristics of strongly correlated insulators.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers JP22K03477 and JP25K07160 and World Premier International Research Center Initiative (WPI), MEXT, Japan. Numerical calculations were partly performed on the Numerical Materials Simulator at National Institute for Materials Science.
@@ A.1 Particle–hole symmetry
Thus, $A_{X^{\prime}}({\bm{k}},\omega)$ of the electron-doped state $|X^{\prime}\rangle$ is obtained as $A_{X}({\bm{\pi}}-{\bm{k}},-\omega)$ of the hole-doped state $|X\rangle$, and vice versa. According to Eqs. (163)–(168),
@@ A.2 Spectral function of nonzero-$\eta$ states
In systems with $\eta$-SU(2) symmetry (Sec. III.1.2), by applying $({\hat{\eta}}^{+})^{m}$ to a $2m$-hole-doped state ($\eta^{z}=-m$) or by applying $({\hat{\eta}}^{-})^{m}$ to a $2m$-electron-doped state ($\eta^{z}=m$), an undoped state ($\eta^{z}=0$) can be obtained. The spectral function of an undoped state of $\eta=n(>0)$ for $\mu=H=0$, $A_{n,0}({\bm{k}},\omega)$, can be obtained from those of the $2m$-electron-doped and $2m$-hole-doped states of $\eta=n$ for $\mu=H=0$, $A_{n,\pm m}({\bm{k}},\omega)$, as
@@
for $n\leq m\leq n$, where $A_{n,m}({\bm{k}},\omega)$ denotes the spectral function of the state of $\eta=n$ and $\eta^{z}=m$ in an $\eta$-SU(2) multiplet. Equation (A.2) is obtained as follows.
@@
where $(T^{\frac{1}{2};{\bm{k}},\sigma}_{\frac{1}{2}},T^{\frac{1}{2};{\bm{k}},\sigma}_{-\frac{1}{2}})=(c^{\dagger}_{{\bm{k}},\sigma},2s^{z}c_{{\bm{\pi}}-{\bm{k}},{\bar{\sigma}}})$, $Z_{n,m}=\langle n,m|n,m\rangle=Z_{n}$, and $i\in(l)$ means that $|i\rangle$ has $\eta=l$. The quantities with subscript $i$ indicate those for $|i\rangle$, and the excitation energy of $|i\rangle$ from $|n,m\rangle$ is denoted as $\epsilon_{i}$.
@@
Equation (A.2) means that $A^{c}_{n,m}({\bm{k}},\omega)+A^{c}_{n,m}({\bm{\pi}}-{\bm{k}},-\omega)$ does not depend on $m$ at $\mu=0$:

### v2

@@ VI.3 Spin-charge separation of interacting metals on a chain
Spin-charge separation means that the lowest spin-excitation energy differs from the lowest charge-excitation energy. In interacting metals on a chain, it is known that the velocity of spin excitation generally differs from that of charge excitation; spin-charge separation occurs in the energy scale of $\mathcal{O}(\frac{1}{L})$, where $L$ denotes the number of sites on a chain Essler ; HaldaneTLL ; TomonagaTLL ; LuttingerTLL ; MattisLiebTLL ; Giamarchi . This characteristic is contrasted with that of a Fermi liquid LandauFL , where spin and charge excitations are obtained as particle-hole excitations of electronic quasiparticles. In a Fermi liquid, the lowest spin-excitation energy is the same as the lowest charge-excitation energy, and spin-charge separation does not occur.
@@ VI.4 Spin-charge separation of strongly correlated insulators
VI.5 Relation between doping-induced in-gap electronic modes and spin-charge separation of strongly correlated insulators
According to the quantum-number analysis (selection rules) for doping-induced modes in Sec. IV.2, the dispersion relations of the doping-induced in-gap modes are expressed [Eqs. (45) and (50)] as
@@
which implies that the emergent in-gap electronic modes exhibit the spin-excitation dispersion relation ($\pm e^{\rm spin}_{\pm{\bm{k}}}$) shifted by the Fermi momentum ($\mp{\bm{k}}_{\rm F}^{\mp}$) from the band edges ($\mu_{\mp}$).
@@
Thus, the doping-induced in-gap modes reflect spin-charge separation of strongly correlated insulators KohnoMottT ; KohnoTinduced ; KohnoRPP ; Kohno1DHub ; Kohno2DHub ; Kohno1DtJ ; Kohno2DtJ ; KohnoDIS ; KohnoHubLadder ; KohnoSpin ; KohnoAF ; KohnoGW ; KohnoKLM ; KohnottpHub ; KohnottpJ ; these do not appear in conventional band insulators without spin-charge separation.
The emergence of in-gap modes in the electronic excitation from the lowest-energy state of a nonzero value of $\eta$ at half filling ($\eta^{z}=0$) [Sec. IV.6; Figs. 6(c), 6(f), 7(c), 8(c), 8(f), 8(i), 8(l), 8(o), 8(r), 8(u), 8(x), 9(c), 9(f), 9(i), and 9(l)] also reflects spin-charge separation.
VI.6 Relation between magnetization-induced in-gap electronic modes and spin-charge separation of strongly correlated insulators
@@ VI.6.1 Emergence of magnetization-induced modes within the band gap in spin-gapped Mott and Kondo insulators
Similarly, according to the quantum-number analysis (selection rules) for magnetization-induced modes in Sec. V.2, the dispersion relations of the magnetization-induced in-gap modes are expressed [Eqs. (80) and (81)] as
@@
The spin gap at $H=0$ is denoted by $\Delta_{\rm spin}$.
@@
From the above consideration, the formation of the spin mode (triplon mode) separated from the particle-hole continuum can be considered the key characteristic for the emergence of electronic modes with magnetization. As shown in Figs. 4 and 5, the spin mode (triplon mode) appears below the particle-hole continuum, and the charge mode appears above the particle-hole continuum in spin-gapped Mott and Kondo insulators. Because the lowest excitation energy of the spin mode is lower than that of the particle-hole continuum, which is equal to the lowest charge-excitation energy, the formation of the spin mode below the particle-hole continuum is a manifestation of spin-charge separation in spin-gapped Mott and Kondo insulators.
Thus, the emergence of magnetization-induced modes within the band gap reflects spin-charge separation, which is realized as the formation of the spin mode below the particle-hole continuum, in spin-gapped Mott and Kondo insulators.
The emergence of in-gap modes in the electronic excitation from the lowest-energy state of a nonzero value of $S$ without magnetization at half filling [Sec. V.6; Figs. 11(c), 11(f), 11(i), 11(l), 11(o), 11(r), 11(u), 11(x), 12(c), 12(f), 12(i), and 12(l)] also reflects spin-charge separation.
@@ VI.6.4 Strongly correlated insulators of attractively interacting electrons
In the insulators of attractively interacting electrons ($U<0$) which are obtained by applying the Shiba transformation (Sec. V.5) to spin-gapped Mott and Kondo insulators with repulsive interactions ($U>0$), because the roles of spin and charge are interchanged by the Shiba transformation [Eqs. (104) and (105)], the charge mode appears below the particle-hole continuum. In this case, the doping-induced modes appear within the band gap, reflecting spin-charge separation (the lowest charge-excitation energy is lower than the lowest spin-excitation energy, which is opposite to the repulsive case).
@@ VI.7 Implications
Although various interpretations of electronic states in strongly correlated systems have been considered KohnoMottT ; KohnoTinduced ; KohnoRPP ; Kohno1DHub ; Kohno2DHub ; Kohno1DtJ ; Kohno2DtJ ; KohnoDIS ; KohnoHubLadder ; KohnoSpin ; KohnoAF ; KohnoGW ; KohnoKLM ; KohnottpHub ; KohnottpJ ; Eskes ; DagottoDOS ; CPTPRL ; CPTPRB ; CPTHanke ; TsunetsuguRMP ; localRungHubLadder ; lightPump1dHub ; pumpProbe1dEHM ; photoSubbands ; FKFloquetDMFT ; noneqDMFTRMP ; DCAkw1dHub ; entropyCoolingDMFT ; etaPairingSuper ; nonthermalSuper ; KanekoEtaPairing ; etaPairingHubLad ; Essler ; HaldaneTLL ; TomonagaTLL ; LuttingerTLL ; MattisLiebTLL ; Giamarchi ; DDMRGAkw ; HubbardI ; LandauFL ; BrinkmanRice ; SlaterAF ; PennAF ; ImadaRMP ; SakaiImadaPRL ; SakaiImadaPRB ; SakaiJPSJ ; ImadaCofermionPRL ; ImadaCofermionPRB ; PhillipsRMP ; PhillipsRPP ; EderOhta2DHub ; EderOhtaIPES ; PreussQP ; GroberQMC ; NoceraDMRT_T ; Akw2dtJ_T , the view presented in this paper along with Refs. [KohnoTinduced, ; KohnoMottT, ; KohnoRPP, ; Kohno1DHub, ; Kohno2DHub, ; Kohno1DtJ, ; Kohno2DtJ, ; KohnoDIS, ; KohnoHubLadder, ; KohnoSpin, ; KohnoAF, ; KohnoGW, ; KohnoKLM, ; KohnottpHub, ; KohnottpJ, ] can consistently explain unconventional features of the electronic states induced not only by temperature increase, doping, and magnetization but also by spin and charge perturbations that cause nonequilibrium states around Mott and Kondo insulators (Secs. IV and V). Furthermore, this view can clarify the differences from the conventional features of the temperature-, doping-, and magnetization-independent band structure of noninteracting band insulators (Sec. VI).
According to this view, electronic modes, whose dispersion relations can be expressed as combinations of the electronic and either the spin or the charge dispersion relations of unperturbed systems, can be induced by various spin and charge perturbations as well as by temperature increase in strongly correlated insulators KohnoMottT ; KohnoTinduced ; KohnoRPP ; Kohno1DHub ; Kohno2DHub ; KohnoDIS ; Kohno1DtJ ; Kohno2DtJ ; KohnoSpin ; KohnoAF ; KohnoGW ; KohnottpHub ; KohnottpJ ; KohnoHubLadder ; KohnoKLM . This characteristic can be applied to band-structure engineering, which makes it possible to increase the number of electronic bands, control the band gap, and even change the insulating band structure to a metallic one at half filling by appropriately tuning and selecting perturbations and strongly correlated materials.
@@ VII Summary
Building upon the understanding of doping- or temperature-induced electronic modes exhibiting momentum-shifted magnetic dispersion relations, this study generalizes the underlying mechanism and demonstrates that various spin and charge perturbations can induce electronic modes and alter the band structure in strongly correlated insulators, provided that a macroscopic number $\mathcal{O}(N_{\rm u})$ of spins or charges are excited. The origins and dispersion relations of these emergent electronic modes, particularly why and how the dispersion relations depend on the momentum and energy of the perturbations, are elucidated by investigating the selection rules and using the Bethe ansatz and the effective theory for weak inter-unit-cell hopping. The validity and generality of the theoretical results across different models and spatial dimensions are verified by numerical calculations for the 1D and 2D HMs, PAMs, KLMs, and ladder and bilayer HMs.
@@
For charge perturbations, not only the chemical-potential shift that causes hole or electron doping but also the perturbation that generates the lowest-energy state of a nonzero value ($\bar{\eta}$}) of $\eta$ at half filling ($N_{\rm h}=0$) can induce electronic modes. The induced electronic modes at half filling exhibit the same dispersion relations as those of the doping-induced modes at $|N_{\rm h}|=2{\bar{\eta}}$ for $\mu=0$, reflecting $\eta$-SU(2) symmetry. In addition, the charge fluctuation that leads to effectively a single dominant state at a given momentum in the charge excitation can induce electronic modes. The charge perturbations at half filling can be realized by photon (light) radiation.
@@
For spin perturbations, the magnetic field that magnetizes spin-gapped Mott and Kondo insulators can induce electronic modes within the band gap. The spin perturbation that generates the lowest-energy state of a nonzero value ($\bar{S}$}) of $S$ without magnetization ($S^{z}=0$) can also induce electronic modes, which exhibit the same dispersion relations as those of the magnetization-induced modes at $|S^{z}|={\bar{S}}$ for $H=0$, reflecting spin-SU(2) symmetry. In addition, the spin fluctuation that leads to effectively a single dominant state at a given momentum in the spin excitation can induce electronic modes. The spin perturbations without a magnetic field can be realized through neutron or photon radiation.
@@
The dispersion relations of the emergent electronic modes can be simply expressed as combinations of the electronic and either the spin or the charge dispersion relations of unperturbed systems because the emergent modes exhibit the dispersion relations of unperturbed systems shifted by the momentum and energy (i.e., a point in the dispersion relations) of the perturbations. The results can be used to interpret changes in the band structure caused by perturbations, including the emergence of electronic states in nonequilibrium systems. Furthermore, the energy and momentum regimes of the emergent electronic states can be controlled by tuning and selecting perturbations and strongly correlated materials based on these results. This approach can be applied to the basic principle of band-structure engineering for strong-correlation electronics, utilizing the unconventional characteristics of strongly correlated insulators.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grants No. JP22K03477 and No. JP25K07160 and World Premier International Research Center Initiative (WPI), MEXT, Japan. Numerical calculations were partly performed on the Numerical Materials Simulator at NIMS.
Data availability
The data that support the findings of this article are openly available dataAvailability .
@@ A.1 Particle–hole symmetry
Thus, $A_{X^{\prime}}({\bm{k}},\omega)$ of the electron-doped eigenstate $|X^{\prime}\rangle$ is obtained as $A_{X}({\bm{\pi}}-{\bm{k}},-\omega)$ of the hole-doped eigenstate $|X\rangle$, and vice versa. According to Eqs. (165)–(170),
@@ A.2 Spectral function of nonzero-$\eta$ states
In the HM, symmetric PAM, and KLM on a bipartite lattice, by applying $({\hat{\eta}}^{+})^{m}$ to a $2m$-hole-doped state ($\eta^{z}=-m$) or by applying $({\hat{\eta}}^{-})^{m}$ to a $2m$-electron-doped state ($\eta^{z}=m$), an undoped state ($\eta^{z}=0$) can be obtained (Sec. III.1.2). The spectral function of an undoped eigenstate of the Hamiltonian with $\eta=n(>0)$ for $\mu=H=0$, $A_{n,0}({\bm{k}},\omega)$, can be obtained from those of the $2m$-electron-doped and $2m$-hole-doped eigenstates with $\eta=n$ for $\mu=H=0$, $A_{n,\pm m}({\bm{k}},\omega)$, as
@@
for $n\leq m\leq n$, where $A_{n,m}({\bm{k}},\omega)$ denotes the spectral function of the state with $\eta=n$ and $\eta^{z}=m$ in an $\eta$-SU(2) multiplet. Equation (A.2) is obtained as follows:
@@
where $(T^{\frac{1}{2};{\bm{k}},\sigma}_{\frac{1}{2}},T^{\frac{1}{2};{\bm{k}},\sigma}_{-\frac{1}{2}})=(c^{\dagger}_{{\bm{k}},\sigma},2s^{z}c_{{\bm{\pi}}-{\bm{k}},{\bar{\sigma}}})$, $Z_{n,m}=\langle n,m|n,m\rangle=Z_{n}$, and $i\in(l)$ means that $|i\rangle$ has $\eta=l$. The quantities with subscript $i$ indicate those for $|i\rangle$, and the excitation energy of $|i\rangle$ from $|n,m\rangle$ is denoted by $\epsilon_{i}$.
@@
Equation (A.2) means that $A^{c}_{n,m}({\bm{k}},\omega)+A^{c}_{n,m}({\bm{\pi}}-{\bm{k}},-\omega)$ does not depend on $m$ at $\mu=0$,